what is the probability of choosing exactly one woman among 2 women and 4 men forming 3 commute.
Answer by stanbon(75887)

what is the probability of choosing exactly one woman among 2 women and 4 men forming 3 committee.
---
Ways to pick a woman: 2
Ways to pick 2 men: 4C2 = 6
----
Ways to pick 3 persons randomly: 6C3 = (6*5*4)/(1*2*3) = 20
-----
P(1 woman on 3 person committee) = (2*6)/20 = 12/20 = 3/5
=====================
Cheers,
Stan H.